OpenStudy (anonymous):

if nell works 3/4 of an hour on monday and 6/10 of an hour on tusday how much time will she work altoghter

OpenStudy (anonymous):

i think its 4

OpenStudy (anthonyn2121):

(3/4)60=45 (6/10)60=36 45+36=81 minutes

OpenStudy (raffle_snaffle):

which is 1 hour and 21min

OpenStudy (anonymous):

joe was makeing mixed nuts . he used 2 /2/4 pounds of pecans and 3/1/5 pounds of cashews .how many pounds of mixed nuts did joe make

OpenStudy (anonymous):

do you guys know the answer yes or no

OpenStudy (anthonyn2121):

What does 2/2/4 mean? Is it 2 and a half pounds?